# refrences.bib

### v1

@@
Abstract
@@
Keywords: Bipolar fuzzy relational equations, Archimedean t-norms, Max-Archimedean compositions, Global optimization, Linear optimization.
@@ 1 Introduction
The theory of fuzzy relational equations (FRE) refers to a generalized version of Boolean relation equations. FRE was originally developed by Sanchez as a formalized model for non-precise concepts [ref_39]. Many applications such as medical diagnosis and treatment are based on fuzzy set theory due to its several properties which make it suitable for formulating uncertain information. Hence, FRE theory was originally applied to medical diagnosis [ref_39]. Considering the sets and various operations, Pedrycz identified two different ways to extend FRE [ref_35]. Throughout the years, FRE has been utilized in a variety of applications in engineering. Upon applying the inference rules and identifying the relevant outcomes, the problem of deriving precedence is eased and mathematically reduced to solving an FRE [ref_33]. The fact remains, however, that a large number of issues can potentially be viewed as FRE issues [ref_34]. Due to the fact that FRE holds a number of important applications in various practical fields, many authors have focused their attention on FRE theoretical research, such as its resolution approach and specific optimization problems.
[…]
@@
Recent studies have demonstrated that the optimization problem involving FRE and FRI is one of the most fascinating and engaging research topics of its kind [ref_01, ref_10, ref_13, ref_14, ref_16, ref_17, ref_18, ref_19, ref_20, ref_21, ref_22, ref_24, ref_30, ref_42, ref_45, ref_51]. Due to the effectiveness of integer linear programming (ILP) for this class of problem, as a commonly used method, researchers have designed various methods based on translating the main problem into an ILP problem. This mapping enables well-established techniques to efficiently solve the problem. Alternatively, other algorithms emphasize the resolution of the feasible region and outline some necessary and sufficient conditions for simplification and optimization processes. The majority of methods in this category leverage analytical results, which are mostly presented by Sanchez [ref_40] and Pedrycz [ref_36]. As an example, Fang and Li mapped a linear optimization problem involving FRE with max-min composition into an integer programming problem and solved it by a branch-and-bound method using a jump-tracking technique [ref_11]. In order to improve the methods that were used by Fang and Li, Wu et al. focused on decreasing the search domain, and based on a necessary condition, they derived three rules for the simplification process [ref_46]. In Chang and Shieh’s work [ref_01], by analyzing the linear optimization problem subject to fuzzy max–min relation equations, they proposed some novel theoretical results that yielded an enhanced upper bound on the optimal objective value, determining some rules for simplifying the problem and suggesting a rule for reducing the solution tree. As an application of optimizing the linear objective with max-min composition, the author in [ref_23], implemented a three-tier framework for the streaming media provider to achieve the minimum cost while meeting the requirements. In other approaches, such as Loetamonphong and Fang, they separated negative from non-negative coefficients in the objective function, dividing the objective function into two subproblems, and by combining the optimal solutions from the two subproblems, an optimal solution was found [ref_31]. In addition, many authors have studied modifications of the linear optimization problem subject to FRE that were defined with other compositions rather than max-min or max-product, like max-average composition [ref_45] or max-t-norm composition [ref_16, ref_17, ref_19, ref_21, ref_24, ref_42]. Accordingly, in order to address the linear optimization problem subjected to a system of sup-t equations, Li and Fang reduced the problem to a 0-1 integer optimization problem [ref_24]. Authors in [ref_21] outline a very method that deals with linear optimization problems involving fuzzy relation equations with constraints defined with the max-Archimedean t-norm. It is worth mentioning that a study reported in [ref_42] examined the problem in equal detail with the continuous Archimedean t-norm that employed the covering problem instead of the branch-and-bound methods.
@@
Over the last few years, researchers in the field of FRE have often investigated the generalization form of linear programming applied to the system of fuzzy relations. They have explored it further by considering the composite operations used in FRE, fuzzy relations used in constraint definition, their impact on the objective function of the problems, and other approaches [ref_05, ref_07, ref_12, ref_19, ref_27, ref_30, ref_48, ref_52]. In this context, Wu et al. outline an efficient method for the optimization of linear fractional programming problems involving FRE with a max-Archimedean t-norm composition [ref_48]. Accordingly, Dempe and Ruziyeva aimed at extending the fuzzy linear optimization problem by taking fuzzy coefficients into account [ref_05]. A further study was conducted by Dubey et al. [ref_07] regarding linear programming problems that utilize interval uncertainty modeled in terms of intuitionist fuzzy sets. As reported by Yang [ref_52], the optimal solution to the minimization of a linear objective function examined under the given constraints is referred to as $\sum_{j=1}^{n}min\left\{a_{ij},x_{j}\right\}\geq b_{i}$ for $i=1,\dots,m$. As well, in [ref_51], the writers discussed the latticized linear programming problem subject to FRI that the constraint is defined as max-product with application to the optimization management model for wireless communication emission base stations. The objective function for the latticized linear programming problem was defined as $z(x)=\min_{j=1}^{n}\{x_{j}\}$ within the feasible region $X(A,b)=\left\{x\in[0,1]^{n}:A\circ x\geq b\right\}$ that $\circ$” refers to the fuzzy max-product composition, for which an algorithm was proposed by considering the resolution of the feasible region.
@@
A bipolar fuzzy relation equation is an intriguing generalization of the fuzzy relation equation. Information processing and preference in humans can be viewed as the main example of bipolarity [ref_08]. It should be noted that Dubois and Prade have reported a study about the relation between asymmetric bipolar and possibility theory [ref_09]. As discussed, possibility theory provides a suitable framework for dealing with bipolar representations, as it allows differentiation between negative and positive information in preference modeling [ref_08, ref_09]. An examination of a linear optimization problem under bipolar FRE was also the main interest of various authors where FRE was defined with max-min (with application to evaluate public awareness of the products of a supplier) [ref_12, ref_26], max-product [ref_04, ref_58] and max-Lukasiewicz composition [ref_27, ref_30, ref_50, ref_53]. The first attempt to address the concept of bipolar FRE was published in [ref_12] which applied max-min composition where constraints are defined as $max_{j=1}^{n}\left\{max\left\{min\left\{a_{ij}^{+},x_{j}\right\},min\left\{a_{ij}^{-},1-x_{j}\right\}\right\}\right\}$ for $i=1,\dots,m$, where $a_{ij}^{+},a_{ij}^{-},x_{j}\in[0,1]$. Likewise, the writers of [ref_27] outlined a linear optimization problem based upon a bipolar FRE system defined as $X\left(A^{+},A^{-},b\right)=\left\{x\in[0,1]^{m}:x\circ A^{+}\vee\tilde{x}\circ A^{-}=b\right\}$ where $\tilde{x}_{i}=1-x_{i}$ for each element of $\tilde{x}=\left(\tilde{x}_{i}\right)_{1\times m}$ and the notations $\vee$” and $\circ$” respectively refer to max operation and the max-Lukasiewicz composition. The authors resolved the problem by translating the original problem into a 0-1 integer linear problem. Alternatively, the discussed problem was addressed by an analytical method in [ref_30] that is motivated by the resolution and some structural properties of the feasible region (by applying a necessary condition for the determination of an optimal solution and a simplification process to reduce the problem). There is, however, no mention of the resolution method used in [ref_50] for determining the complete solution set of the bipolar max-Lukasiewicz FRE. As a result of Yang’s research on bipolar max-Lukasiewicz FRE [ref_50], it has been demonstrated that the complete solution set of the system is completely specified by a finite number of conservative bipolar paths. In addition, the authors in [ref_60] studied a more general class of bipolar FREs defined by max strict t-norms. They categorized the constraints of the problem into four groups, each requiring a specific approach to determine the corresponding feasible solution set. Moreover, unlike the bipolar path approach [ref_50], the authors demonstrated that modified ordinary FRE paths could be used to solve the problem, which is achieved by combining the FRE paths specified in the four groups of the constraints [ref_60].
@@
It is worth mentioning that since the objective function can be transformed into an increasing function using the change of variables [ref_12], it can be assumed without loss of generality that $c_{j}\geq 0$, $\forall i\in\mathscr{I}$. In the current method, unlike the method presented in [ref_60], to find a global optimum, it is not necessary to obtain all local optimal solutions. Moreover, by using the current method, the constraints do not need to be classified (and therefore, it is unnecessary to apply different approaches to check the feasibility of constraints in different groups and the optimality of solutions in their feasible regions). It is proved that the feasible region of Problem (1) can be determined as a finite number of compact sets, which are obtained by functions called admissible functions in this paper. After deriving some basic properties of Archimedean bipolar fuzzy equations, seven simplification techniques are presented which are applied at the beginning of the algorithm to reduce the problem size; four techniques are based on eliminating redundant constraints, and three additional techniques eliminate a set of feasible solutions that are not optimal. Subsequently, after applying the simplification techniques, the proposed algorithm takes advantage of two accelerating strategies to find an optimal solution. The first strategy removes some candidate solutions that are not optimal (by reducing the domain search to a subset of admissible functions called modified functions in this paper), and the second strategy (a branch and bound method), removes some branches that do not lead to an optimal solution.
@@ 2 Preliminary definitions and properties
It should be noted that if $S\neq\varnothing$ in Definition 1 , then both $l=\inf S$ and $u=\sup S$ exist. This consequence is actually resulted from the fact that $S\subseteq[0,1]$ and the least-upper-bound property of the real numbers set.
Lemma 1.
(a) If $x_{0}\in S$, then $x\in I$ for each $x\in\left[0,x_{0}\right]$. (b) If $a<b$, then $I=[0,1]$ and $S=\varnothing$.
@@
    (a) Since $x_{0}\in S$, then from Definition 1 we have $\varphi\left(a,x_{0}\right)=b$. On the other hand, from the monotonicity property of t-norms, $\varphi(a,x)\leq\varphi\left(a,x_{0}\right)$, $\forall x\in\left[0,x_{0}\right]$. So, $\varphi(a,x)\leq b$, $\forall x\in\left[0,x_{0}\right]$, which together with Definition 1 imply $x\in I$, $\forall x\in\left[0,x_{0}\right]$. (b) The result follows from the identity and monotonicity laws of t-norms and the inequalities $\varphi(a,x)\leq\varphi(a,1)=a<b$, $\forall x\in[0,1]$. ∎
Lemma 2.
[ref_61] A function $\varphi:[0,1]^{2}\rightarrow[0,1]$ is a continuous Archimedean t-norm if and only if there exist a strictly decreasing and continuous function $f_{\varphi}:[0,1]\rightarrow[0,+\infty)$ with $f_{\varphi}(1)=0$ such that
@@
Moreover, representation (3) is unique up to a positive multiplicative constant.
We say that $\varphi$ is generated by $f_{\varphi}$ if $\varphi$ has representation (3). In this case, $f_{\varphi}$ is said to be an additive generator of $\varphi$. For example, an additive generator of Product tnorm is given by $f_{P}(x)=-\operatorname{Ln}(x)$. Then, $f_{P}^{(-1)}(x)=e^{-x}$. Also, Lukasiewicz t-norm is generated by $f_{L}(x)=1-x$ with $f_{L}^{(-1)}(x)=\max\{1-x,0\}$. Table A1 (see Appendix A) lists some frequently used continuous Archimedean t-norms (including strict and nilpotent $\mathrm{t}$-norms). For each $\mathrm{t}$-norm in Table A1, its additive generator $f_{\varphi}$ and pseudoinverse $f_{\varphi}^{(-1)}$ discussed in Lemma 2 are presented in Table A2.
Remark 1.
Suppose that $a\in[0,1]$ and $0<b\leq 1$. Based on Lemma 2, the equation $\varphi(a,x)=b$ can be rewritten as follows:
@@
Theorem 1.
Suppose that $\varphi$ is a continuous Archimedean t-norm, $f_{\varphi}$ is an additive generator of $\varphi$ and $u=f_{\varphi}^{-1}\left(f_{\varphi}(b)-f_{\varphi}(a)\right)$, where $a\geq b>0$. (a) $S=\{u\}$ and $I=[0,u]$. (b) If $b<a<1$, then $b<u<1$. (c) If $b<a=1$, then $u=b$. (d) If $a=b$, then $u=1$.
@@
    (a) From Remark 1 (relation (5)), $\varphi(a,x)=b$ if and only if $f_{\varphi}^{-1}\left(f_{\varphi}(a)+f_{\varphi}(x)\right)=b$. Now, since $f_{\varphi}$ and $f_{\varphi}^{-1}$ are strictly decreasing continuous functions, it is concluded that $f_{\varphi}^{-1}\left(f_{\varphi}(a)+f_{\varphi}(x)\right)=b$ if and only if $x=u=f_{\varphi}^{-1}\left(f_{\varphi}(b)-f_{\varphi}(a)\right)$. Hence, $S=\{u\}$, which together with Lemma 1 (a) imply $I=[0,u]$. (b) From Part (a), we have $S=\{u\}$. Since $\varphi(a,1)=a>b$, then $1\notin S$, which implies $u<1$. On the other hand, for each $x\in[0,b)$, $\varphi(a,x)\leq\varphi(1,x)=x<b$, that means $x\notin S$, $\forall x\in[0,b)$. Moreover, $b\notin S$; because otherwise, if $\varphi(a,b)=b$, then the associativity property of t-norms implies $\varphi(a,\varphi(a,b))=\varphi(\varphi(a,a),b)=b$, and by repeating the argument we have $\varphi\left(\varphi^{(n)}(a),b\right)=b$ ($\forall n\in\mathbb{R}^{n}$), where $\varphi^{(n)}(a)$ denotes $\varphi(\underbrace{a,\ldots,a}_{n\text{-times }})$. Therefore,
@@
Theorem 2.
Suppose that $\varphi$ is a continuous Archimedean t-norm, $f_{\varphi}$ is an additive generator of $\varphi$ and $u=f_{\varphi}^{-1}\left(f_{\varphi}(b)-f_{\varphi}(a)\right)$, where $a\geq b=0$. (a) If $a=b=0$, then $S=I=[0,u]$ and $u=1$. (b) If $a>b=0$, then $S=I=[0,u]$ and $10\leq u<$ 1.
@@
    (a) If $a=b=0$, then the equation $\varphi(a,x)=b$ is converted to $\varphi(0,x)=0$, which is satisfied for each $x\in[0,1]$. Hence, $S=[0,1]$ and therefore $I=[0,1]$ from Lemma 1 (a). Also, $u=f_{\varphi}^{-1}\left(f_{\varphi}(b)-f_{\varphi}(a)\right)=f_{\varphi}^{-1}\left(f_{\varphi}(0)-f_{\varphi}(0)\right)=f_{\varphi}^{-1}(0)=1$. (b) In this case, from relation (6), the equation $\varphi(a,x)=0$ is converted to $f_{\varphi}^{(-1)}\left(f_{\varphi}(a)+f_{\varphi}(x)\right)=0$, which together with (4) imply
@@
    So, $\varphi(a,u)=0$ which together with (10) imply $\varphi(a,x)=0$, $\forall x\in[0,u]$; that is, $[0,u]\subseteq S$. In order to prove $S=[0,u]$, we shall show that $x\notin S$ for each $x>u$. By contradiction, suppose that $\varphi(a,x)=0$ (i.e., $x\in S$ ) and $x>u=f_{\varphi}^{-1}\left(f_{\varphi}(0)-f_{\varphi}(a)\right)$. Since $\varphi(a,x)=0$, then $x$ satisfies (9). On the other hand, since $x>f_{\varphi}^{-1}\left(f_{\varphi}(0)-f_{\varphi}(a)\right)$, it is concluded that $f_{\varphi}(x)<f_{\varphi}(0)-f_{\varphi}(a)$, that contradicts (9). Consequently, $S=[0,u]$ and therefore $I=[0,u]$ from Lemma 1 (a). Moreover, since $\varphi(a,1)=a>b$, it follows that $1\notin S=[0,u]$, i.e., $u<1$. ∎
From Lemma 1 (b), Theorem 1, and Theorem 2, we can obtain a complete characterization of $S$ when $\varphi$ is a continuous Archimedean t-norm (i.e., $\varphi$ is either continuous strict or continuous nilpotent). These results are summarized in the following corollary.
Corollary 1.
Suppose that $\varphi$ is a continuous Archimedean t-norm, $f_{\varphi}$ is an additive generator of $\varphi$ and $u=f_{\varphi}^{-1}\left(f_{\varphi}(b)-f_{\varphi}(a)\right)$. Then,

### v2

@@
[…]
On the resolution and linear optimization problems subject to a system of bipolar fuzzy relational equalities defined with continuous Archimedean t-norms 11affiliationtext: Faculty of Engineering Science, College of Engineering, University of Tehran, P.O.Box 11365-4563, Tehran, Iran. 22affiliationtext: Department of Engineering Science, College of Engineering, University of Tehran, Tehran, Iran. 33affiliationtext: Department of Electrical and Computer Engineering, University of Alberta, Edmonton, AB T6G 2R3, Canada 44affiliationtext: Institute of Systems Engineering, Macau University of Science and Technology, Taipa 999078, Macau SAR, China 55affiliationtext: Research Center of Performance and Productivity Analysis, Istinye University, Istanbul, Türkiye
[…]
@@
[…]
Bipolar fuzzy relational equations, Archimedean t-norms, Max-Archimedean compositions, Global optimization, Linear optimization.
[…]
@@ 1 Introduction
The theory of fuzzy relational equations (FRE) refers to a generalized version of Boolean relation equations. FRE was originally developed by Sanchez as a formalized model for non-precise concepts [39]. Many applications such as medical diagnosis and treatment are based on fuzzy set theory due to its several properties which make it suitable for formulating uncertain information. Hence, FRE theory was originally applied to medical diagnosis [39]. Considering the sets and various operations, Pedrycz identified two different ways to extend FRE [37]. Throughout the years, FRE has been utilized in a variety of applications in engineering. Upon applying the inference rules and identifying the relevant outcomes, the problem of deriving precedence is eased and mathematically reduced to solving an FRE [35]. The fact remains, however, that a large number of issues can potentially be viewed as FRE issues [38]. Due to the fact that FRE holds a number of important applications in various practical fields, many authors have focused their attention on FRE theoretical research, such as its resolution approach and specific optimization problems.
In dealing with FRE problems, it is imperative to identify the solvability of the problem and determine the set of solutions. Based on the study reported in [36], the authors concluded that the solution set for FRE with a continuous max-t-norm composition (when it is not empty), could be a non-convex set, which is entirely determined with only one maximum solution and a finite number of minimal solutions. It faces two significant bottlenecks when it comes to tackling FRE-related problems. The first is the non-convexity property of solution sets. This property significantly increases the complexity of problems, especially in the case of optimization problems involving FREs. The second bottleneck lies in efficiently identifying minimal solutions. As described in Chen and Wang’s study [2], they proposed an algorithm for extracting all minimal solutions and concluded that there might not be a polynomial-time algorithm for finding all minimal solutions of FRE (with max-min composition). Furthermore, Markovskii demonstrated a close connection between solving max-product FREs and the covering problem, a well-known NP-hard problem [34]. Moreover, this result extends beyond the minimum and product t-norm operators. Indeed, it generalizes to a wider class of t-norm functions [3, 31, 30]. In the past few decades, many researchers have attempted to investigate the solvability of FRE through different max-t compositions [20, 14, 41, 43, 44]. Building on prior research, several scholars have made significant contributions to the field of fuzzy relational inequalities (FRI). This effort has introduced new concepts, improved time complexity, and refined existing theoretical structures and applications [19, 15, 14, 21, 26, 50, 51]. In a noteworthy contribution to the field, Li and Yang [26] examined FRI using addition-min composition and developed an algorithm to search for minimal solutions. Their work centered on applying FRI to the data transmission mechanism in a BitTorrent-like peer-to-peer file-sharing system. As noted in [19], a mixed fuzzy system was studied by combining two fuzzy relational inequalities $A\varphi x\leq b^{1}$ and $D\varphi x\geq b^{2}$, where $\varphi$ refers to an operator which has (closed) convex solutions.
@@
Recent studies have demonstrated that the optimization problem involving FRE and FRI is one of the most fascinating and engaging research topics of its kind [1, 11, 18, 19, 20, 15, 14, 21, 23, 22, 27, 32, 42, 45, 50, 17]. Due to the effectiveness of integer linear programming (ILP) for this class of problem, as a commonly used method, researchers have designed various methods based on translating the main problem into an ILP problem. This mapping enables well-established techniques to efficiently solve the problem. Alternatively, other algorithms emphasize the resolution of the feasible region and outline some necessary and sufficient conditions for simplification and optimization processes. The majority of methods in this category leverage analytical results, which are mostly presented by Sanchez [40] and Pedrycz [24]. As an example, Fang and Li mapped a linear optimization problem involving FRE with max-min composition into an integer programming problem and solved it by a branch-and-bound method using a jump-tracking technique [12]. In order to improve the methods that were used by Fang and Li, Wu et al. focused on decreasing the search domain, and based on a necessary condition, they derived three rules for the simplification process [46]. In Chang and Shieh’s work [1], by analyzing the linear optimization problem subject to fuzzy max–min relation equations, they proposed some novel theoretical results that yielded an enhanced upper bound on the optimal objective value, determining some rules for simplifying the problem and suggesting a rule for reducing the solution tree. As an application of optimizing the linear objective with max-min composition, the author in [25], implemented a three-tier framework for the streaming media provider to achieve the minimum cost while meeting the requirements. In other approaches, such as Loetamonphong and Fang, they separated negative from non-negative coefficients in the objective function, dividing the objective function into two subproblems, and by combining the optimal solutions from the two subproblems, an optimal solution was found [33]. In addition, many authors have studied modifications of the linear optimization problem subject to FRE that were defined with other compositions rather than max-min or max-product, like max-average composition [45] or max-t-norm composition [20, 14, 23, 27, 42]. Accordingly, in order to address the linear optimization problem subjected to a system of sup-t equations, Li and Fang reduced the problem to a 0-1 integer optimization problem [27]. Authors in [23] outline a very method that deals with linear optimization problems involving fuzzy relation equations with constraints defined with the max-Archimedean t-norm. It is worth mentioning that a study reported in [42] examined the problem in equal detail with the continuous Archimedean t-norm that employed the covering problem instead of the branch-and-bound methods.
@@
Over the last few years, researchers in the field of FRE have often investigated the generalization form of linear programming applied to the system of fuzzy relations. They have explored it further by considering the composite operations used in FRE, fuzzy relations used in constraint definition, their impact on the objective function of the problems, and other approaches [6, 7, 13, 14, 29, 32, 47, 48]. In this context, Wu et al. outline an efficient method for the optimization of linear fractional programming problems involving FRE with a max-Archimedean t-norm composition [47]. Accordingly, Dempe and Ruziyeva aimed at extending the fuzzy linear optimization problem by taking fuzzy coefficients into account [6]. A further study was conducted by Dubey et al. [7] regarding linear programming problems that utilize interval uncertainty modeled in terms of intuitionist fuzzy sets. As reported by Yang [48], the optimal solution to the minimization of a linear objective function examined under the given constraints is referred to as $\sum_{j=1}^{n}min\left\{a_{ij},x_{j}\right\}\geq b_{i}$ for $i=1,\dots,m$. As well, in [50], the writers discussed the latticized linear programming problem subject to FRI that the constraint is defined as max-product with application to the optimization management model for wireless communication emission base stations. The objective function for the latticized linear programming problem was defined as $z(x)=\min_{j=1}^{n}\{x_{j}\}$ within the feasible region $X(A,b)=\left\{x\in[0,1]^{n}:A\circ x\geq b\right\}$ that $\circ$” refers to the fuzzy max-product composition, for which an algorithm was proposed by considering the resolution of the feasible region.
@@
A bipolar fuzzy relation equation is an intriguing generalization of the fuzzy relation equation. Information processing and preference in humans can be viewed as the main example of bipolarity [9]. It should be noted that Dubois and Prade have reported a study about the relation between asymmetric bipolar and possibility theory [10]. As discussed, possibility theory provides a suitable framework for dealing with bipolar representations, as it allows differentiation between negative and positive information in preference modeling [9, 10]. An examination of a linear optimization problem under bipolar FRE was also the main interest of various authors where FRE was defined with max-min (with application to evaluate public awareness of the products of a supplier) [13, 28], max-product [4, 5] and max-Lukasiewicz composition [29, 32, 49, 52]. The first attempt to address the concept of bipolar FRE was published in [13] which applied max-min composition where constraints are defined as $max_{j=1}^{n}\left\{max\left\{min\left\{a_{ij}^{+},x_{j}\right\},min\left\{a_{ij}^{-},1-x_{j}\right\}\right\}\right\}$ for $i=1,\dots,m$, where $a_{ij}^{+},a_{ij}^{-},x_{j}\in[0,1]$. Likewise, the writers of [29] outlined a linear optimization problem based upon a bipolar FRE system defined as $X\left(A^{+},A^{-},b\right)=\left\{x\in[0,1]^{m}:x\circ A^{+}\vee\tilde{x}\circ A^{-}=b\right\}$ where $\tilde{x}_{i}=1-x_{i}$ for each element of $\tilde{x}=\left(\tilde{x}_{i}\right)_{1\times m}$ and the notations $\vee$” and $\circ$” respectively refer to max operation and the max-Lukasiewicz composition. The authors resolved the problem by translating the original problem into a 0-1 integer linear problem. Alternatively, the discussed problem was addressed by an analytical method in [32] that is motivated by the resolution and some structural properties of the feasible region (by applying a necessary condition for the determination of an optimal solution and a simplification process to reduce the problem). There is, however, no mention of the resolution method used in [49] for determining the complete solution set of the bipolar max-Lukasiewicz FRE. As a result of Yang’s research on bipolar max-Lukasiewicz FRE [49], it has been demonstrated that the complete solution set of the system is completely specified by a finite number of conservative bipolar paths. In addition, the authors in [16] studied a more general class of bipolar FREs defined by max strict t-norms. They categorized the constraints of the problem into four groups, each requiring a specific approach to determine the corresponding feasible solution set. Moreover, unlike the bipolar path approach [49], the authors demonstrated that modified ordinary FRE paths could be used to solve the problem, which is achieved by combining the FRE paths specified in the four groups of the constraints [16].
@@
It is worth mentioning that since the objective function can be transformed into an increasing function using the change of variables [13], it can be assumed without loss of generality that $c_{j}\geq 0$, $\forall i\in\mathscr{I}$. In the current method, unlike the method presented in [16], to find a global optimum, it is not necessary to obtain all local optimal solutions. Moreover, by using the current method, the constraints do not need to be classified (and therefore, it is unnecessary to apply different approaches to check the feasibility of constraints in different groups and the optimality of solutions in their feasible regions). It is proved that the feasible region of Problem (1) can be determined as a finite number of compact sets, which are obtained by functions called admissible functions in this paper. After deriving some basic properties of Archimedean bipolar fuzzy equations, seven simplification techniques are presented which are applied at the beginning of the algorithm to reduce the problem size; four techniques are based on eliminating redundant constraints, and three additional techniques eliminate a set of feasible solutions that are not optimal. Subsequently, after applying the simplification techniques, the proposed algorithm takes advantage of two accelerating strategies to find an optimal solution. The first strategy removes some candidate solutions that are not optimal (by reducing the domain search to a subset of admissible functions called modified functions in this paper), and the second strategy (a branch and bound method), removes some branches that do not lead to an optimal solution.
@@ 2 Preliminary definitions and properties
It should be noted that if $S\neq\varnothing$ in Definition 2.1, then both $l=\inf S$ and $u=\sup S$ exist. This consequence is actually resulted from the fact that $S\subseteq[0,1]$ and the least-upper-bound property of the real numbers set.
Lemma 2.2.
Suppose that $\varphi$ is a continuous Archimedean t-norm, (a) If $x_{0}\in S$, then $x\in I$ for each $x\in\left[0,x_{0}\right]$. (b) If $a<b$, then $I=[0,1]$ and $S=\varnothing$.
@@
(a) Since $x_{0}\in S$, then from Definition 2.1 we have $\varphi\left(a,x_{0}\right)=b$. On the other hand, from the monotonicity property of t-norms, $\varphi(a,x)\leq\varphi\left(a,x_{0}\right)$, $\forall x\in\left[0,x_{0}\right]$. So, $\varphi(a,x)\leq b$, $\forall x\in\left[0,x_{0}\right]$, which together with Definition 2.1 imply $x\in I$, $\forall x\in\left[0,x_{0}\right]$. (b) The result follows from the identity and monotonicity laws of t-norms and the inequalities $\varphi(a,x)\leq\varphi(a,1)=a<b$, $\forall x\in[0,1]$. ∎
Lemma 2.3.
[8] A function $\varphi:[0,1]^{2}\rightarrow[0,1]$ is a continuous Archimedean t-norm if and only if there exist a strictly decreasing and continuous function $f_{\varphi}:[0,1]\rightarrow[0,+\infty)$ with $f_{\varphi}(1)=0$ such that
@@
and $f^{-1}_{\varphi}(x)$ is the inverse of $f_{\varphi}$. Moreover, representation (3) is unique up to a positive multiplicative constant. _
We say that $\varphi$ is generated by $f_{\varphi}$ if $\varphi$ has representation (3). In this case, $f_{\varphi}$ is said to be an additive generator of $\varphi$. For example, an additive generator of Product tnorm is given by $f_{P}(x)=-\operatorname{Ln}(x)$. Then, $f_{P}^{(-1)}(x)=e^{-x}$. Also, Lukasiewicz t-norm is generated by $f_{L}(x)=1-x$ with $f_{L}^{(-1)}(x)=\max\{1-x,0\}$. Table A1 (see Appendix A) lists some frequently used continuous Archimedean t-norms (including strict and nilpotent $\mathrm{t}$-norms). For each $\mathrm{t}$-norm in Table A1, its additive generator $f_{\varphi}$ and pseudoinverse $f_{\varphi}^{(-1)}$ discussed in Lemma 2.3 are presented in Table A2.
Remark 2.4.
Suppose that $a\in[0,1]$ and $0<b\leq 1$. Based on Lemma 2.3, the equation $\varphi(a,x)=b$ can be rewritten as follows:
@@
Theorem 2.5.
Suppose that $\varphi$ is a continuous Archimedean t-norm, $f_{\varphi}$ is an additive generator of $\varphi$ and $u=f_{\varphi}^{-1}(f_{\varphi}(b)$- $f_{\varphi}(a))$, where $a\geq b>0$. (a) $S=\{u\}$ and $I=[0,u]$. (b) If $b<a<1$, then $b<u<1$. (c) If $b<a=1$, then $u=b$. (d) If $a=b$, then $u=1$.
@@
(a) Clearly From Remark 2.4, $\varphi(a,u)=f_{\varphi}^{-1}(f_{\varphi}(a)+f_{\varphi}(u))=f_{\varphi}^{-1}(f_{\varphi}(a)+(f_{\varphi}(b)-f_{\varphi}(a)))$. Also relation (5) imply $\varphi(a,x)=b$ if and only if $f_{\varphi}^{-1}\left(f_{\varphi}(a)+f_{\varphi}(x)\right)=b$. Now, since $f_{\varphi}$ and $f_{\varphi}^{-1}$ are strictly decreasing continuous functions, it is concluded that $f_{\varphi}^{-1}\left(f_{\varphi}(a)+f_{\varphi}(x)\right)=b$ if and only if $x=u=f_{\varphi}^{-1}\left(f_{\varphi}(b)-f_{\varphi}(a)\right)$. Hence, $S=\{u\}$, which together with Lemma 2.2 (a) imply $I=[0,u]$. (b) From Part (a), we have $S=\{u\}$. Since $\varphi(a,1)=a>b$, then $1\notin S$, which implies $u<1$. On the other hand, for each $x\in[0,b)$, $\varphi(a,x)\leq\varphi(1,x)=x<b$, that means $x\notin S$, $\forall x\in[0,b)$. Moreover, $b\notin S$; because otherwise, if $\varphi(a,b)=b$, then the associativity property of t-norms implies $\varphi(a,\varphi(a,b))=\varphi(\varphi(a,a),b)=b$, and by repeating the argument we have $\varphi\left(\varphi^{(n)}(a),b\right)=b$ ($\forall n\in\mathbb{R}^{n}$), where $\varphi^{(n)}(a)$ denotes $\varphi(\underbrace{a,\ldots,a}_{n\text{-times }})$. Therefore,
@@
Theorem 2.6.
Suppose that $\varphi$ is a continuous Archimedean t-norm, $f_{\varphi}$ is an additive generator of $\varphi$ and $u=f_{\varphi}^{-1}(f_{\varphi}(b)$- $f_{\varphi}(a))$, where $a\geq b=0$. (a) If $a=b=0$, then $S=I=[0,u]$ and $u=1$. (b) If $a>b=0$, then $S=I=[0,u]$ and $0\leq u<1$.
@@
(a) If $a=b=0$, then the equation $\varphi(a,x)=b$ is converted to $\varphi(0,x)=0$, which is satisfied for each $x\in[0,1]$. Hence, $S=[0,1]$ and therefore $I=[0,1]$ from Lemma 2.2 (a). Also, $u=f_{\varphi}^{-1}\left(f_{\varphi}(b)-f_{\varphi}(a)\right)=f_{\varphi}^{-1}\left(f_{\varphi}(0)-f_{\varphi}(0)\right)=f_{\varphi}^{-1}(0)=1$. (b) In this case, from relation (6), the equation $\varphi(a,x)=0$ is converted to $f_{\varphi}^{(-1)}\left(f_{\varphi}(a)+f_{\varphi}(x)\right)=0$, which together with (4) imply
@@
So, $\varphi(a,u)=0$ which together with (10) imply $\varphi(a,x)=0$, $\forall x\in[0,u]$; that is, $0,u]\subseteq S$. In order to prove $S=[0,u]$, we shall show that $x\notin S$ for each $x>u$. By contradiction, suppose that $\varphi(a,x)=0$ (i.e., $x\in S$ ) and $x>u=f_{\varphi}^{-1}\left(f_{\varphi}(0)-f_{\varphi}(a)\right)$. Since $\varphi(a,x)=0$, then $x$ satisfies (9). On the other hand, since $x>f_{\varphi}^{-1}\left(f_{\varphi}(0)-f_{\varphi}(a)\right)$, it is concluded that $f_{\varphi}(x)<f_{\varphi}(0)-f_{\varphi}(a)$, that contradicts (4). Consequently, $S=[0,u]$ and therefore $I=[0,u]$ from Lemma 2.2 (a). Moreover, since $\varphi(a,1)=a>b$, it follows that $1\notin S=[0,u]$, i.e., $u<1$. ∎
From Lemma 2.2 (b), Theorem 2.5, and Theorem 2.6, we can obtain a complete characterization of $S$ when $\varphi$ is a continuous Archimedean t-norm (i.e., $\varphi$ is either continuous strict or continuous nilpotent). These results are summarized in the following corollary.
Corollary 2.7.
Suppose that $\varphi$ is a continuous Archimedean t-norm, $f_{\varphi}$ is an additive generator of $\varphi$ and $u=f_{\varphi}^{-1}(f_{\varphi}(b)$- $f_{\varphi}(a))$. Then,